3. 10.00 points Consider the following projects: Cash Flows (S) -10,500 -20,500 21,000 35,875 Assume that the projects are mutually exclusive and that the opportunity cost of capital is 10%. a. Calculate the protitability index for each project. (Do not round intermediate calculations. Round your answers to 2 decimal places.) Project Profitability Index b-1. Calculate the protitablity-index using the incremental cash flows. (Do not round intermediate calculations. Round your answer to 2 decimal places.) Prontability-Index b-2. Which project should you choose? Project D O Project BE
Answer #1

a.

PV of cash inflows for project D= 21000 * (1/1.1) = 19090.91

Profitability index for Project D = 19090.91/10500 = 1.82

PV of cash inflows for project e = 35875*(1/1.1) = 32613.64

Profitability index for project e = 32613.64/20500 = 1.59

b-1

Incremental cash flow for.project D = 21000-10500= 10500

Incremental cash flow for project e = 35875-20500= 15375

Profitability index = 15375-10500= 4875

b-2

Based on Incremental cash flow project e should be chosen since it's incremental cash flow is more than project D incremental cash flow

Note: if point a is to be considered based on profitability index project D should be chosen